My question is: Construction elevator is moving downward at 5 m/s. If the worker on the elevator drops his pliers and it hits the ground 10 seconds later, how high was he when he dropped the pliers?
1
Expert's answer
2019-10-22T10:02:45-0400

The distance traveled by pliers


"d=v_it+\frac{gt^2}{2}"

"=5\:\rm m/s\times 10\: s+\frac{9.8\: m/s^2\times (10\:\ s)^2}{2}=540\: m"